# Rough or crumpled: Strong coupling phases of a generalized Kardar-Parisi-Zhang surface

### v1

@@ Abstract
We study a generalized Kardar-Parisi-Zhang (KPZ) equation [D. Jana et al, Phys. Rev. E 109, L032104 (2024)], that sets the paradigm for universality in roughening of growing nonequilibrium surfaces without any conservation laws, but with competing local and nonlocal nonlinear effects. We show that such a generalized KPZ equation in two dimensions can describe a strong coupling rough or a crumpled surface, in addition to a weak coupling phase. The conformation fluctuations of such a rough surface are given by nonuniversal exponents, with orientational long-ranged order and positional short-ranged order, whereas the crumpled phase has positional and orientational short range order. Experimental and theoretical implications of these results are discussed.
The Kardar-Parisi-Zhang (KPZ) equation Kardar et al. (1986); Krug (1997); Barabási and Stanley (1995), originally proposed as a model for growing surfaces without overhangs, provides a paradigm for nonequilibrium phase transitions. It undergoes a roughening transition between a smooth phase, statistically identical to a surface described by the linear Edwards-Wilkinson (EW) equation Edwards and Wilkinson (1982), and a perturbatively inaccessible strong coupling phase Barabási and Stanley (1995), in dimension $d>2$. Perturbative dynamic renormalization group (RG) Barabási and Stanley (1995); Täuber (2014) analysis has been successful in describing this roughening transition in an $\epsilon$-expansion, where $d=2+\epsilon$, with two dimensions (2D) being the lower critical dimension of the model. In $d\leq 2$, there exists only the strong coupling phase. A KPZ surface described by a single-valued height field $h({\bf x},t)$ measured with respect to an arbitrary base plane in the Monge gauge Nelson et al. (1989) displays universal scaling in the long wavelength limit, characterized by $z$ and $\chi$, the dynamic and roughness exponents Barabási and Stanley (1995). These are formally defined through the correlation function
@@
in the scaling limit, where $r=|{\bf x}|$ and $\phi$ is a dimensionless scaling function of its argument Barabási and Stanley (1995). In the smooth phase, $z=2$ and $\chi=1-d/2$ Barabási and Stanley (1995) and at the roughening transition at $d=2+\epsilon,\,\epsilon>0$, $z=2$ and $\chi=0$ Barabási and Stanley (1995); Täuber (2014); Janssen et al. (1999). Failure of RG to capture the strong coupling phase has led to the development of alternative techniques. Notable are the self-consistent mode-coupling theories (MCT) Bouchaud and Cates (1993); Bhattacharjee (1998), which have predicted a variety of results on scaling within specific different calculational approaches. For example Ref. Hwa and Frey (1991) explored the scenario with $d=1$, yielding $z=3/2$, a value known exactly due to the Galilean invariance and a Fluctuation-Dissipation-Theorem Barabási and Stanley (1995). Furthermore, Ref. Doherty et al. (1994) shows that $z$ rises from $z=3/2$ at $d=1$ to 2 around $d_{U}=3.6$, the upper critical dimension of the KPZ equation. In contrast, Ref. Tu (1994) suggests $z<2$ at any finite dimension indicating $d_{U}=\infty$. Furthermore, Refs. Bhattacharjee (1998); Bhattacharjee and Bhattacharyya (2007); Basu and Frey (2004, 2009a) used MCT to predict $\chi=(4-d)/6,\,z=(8+d)/6$ that tend to agree reasonably with the numerical results Gomes-Filho et al. (2021). See also Ref. Colaiori and Moore (2001), which gives $d_{c}=4$ and dynamic exponent $z$ in different dimensions. More recently, Ref. Canet and Moore (2007) revisited the MCT for the KPZ equation and showed that solutions for the MCT equations for the KPZ equation actually have two branches or two distinct universality classes below $d<2$. At $d\geq 2$, there is only one branch that continues to $d_{U}=4$, giving scaling exponents closed to those known from numerical studies.
@@
The dynamics of a KPZ surface depends locally on surface fluctuations, and hence cannot model nonequilibrium surfaces with nonlocal dynamics. Nonlocal effects, however, can be important in wide-ranging systems, including biological growth process Santalla and Ferreira (2018), fast nonlocal transport Aharonov and Rothman (1996) and nonlocal stabiliztion of surfaces Krug and Meakin (1991); Nicoli et al. (2009). Recently, a generalized or “active” KPZ equation (hereafter a-KPZ equation) with nonlocal effects as a conceptual model with competing local and nonlocal nonlinear effects has been proposed Jana et al. (2024)
$\frac{\partial h}{\partial t}$ $=$ $\nu\nabla^{2}h+\frac{\lambda}{2}(\bm{\nabla}h)^{2}$ (2)
$+$ $\lambda_{1}\int d^{d}rQ_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))+\eta.$
@@
Here, $Q_{ij}({\bf r})$ denotes the longitudinal projection operator, which in Fourier space is expressed as $Q_{ij}({\bf k})=k_{i}k_{j}/k^{2}$, with ${\bf k}$ being a Fourier wave vector. Thus $Q_{ij}({\bf r})$, which is the inverse Fourier transform of $Q_{ij}({\bf k})$, is nonlocal in space. Physically, the term $\lambda_{1}\int d^{d}rQ_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))$ contributes to the surface velocity normal to the base plane, $v_{p}=\partial h/\partial t$, and is nonlocal in ${\bm{\nabla}}h$. Furthermore, $\eta({\bf x},t)$ is a zero-mean Gaussian-distributed white noise with a variance $\langle\eta({\bf x},t)\eta({\bf 0},0)\rangle=2D\delta^{d}({\bf x})\delta(t).$ If $\lambda_{1}=0$, Eq. (2) reduces to the usual KPZ equation. As a consequence of the competition between the local and nonlocal nonlinearities, a-KPZ equation (2) shows macroscopic properties dramatically different from the KPZ equation. It can have a stable, sub- or super-logarithmically rough surface in 2D, characterized by nonuniversal scaling exponents. More strikingly and unlike the 2D KPZ equation, it also admits a roughening transition even in 2D, again characterized by nonuniversal exponents. At $d>2$, it shows a roughening transition different from that in the KPZ equation Jana et al. (2024). Equation (2) admits a pseudo-Galilean invariance under the transformation $x_{i}^{\prime}=x_{i}-(\lambda+2\lambda_{1})c_{i}t$ and $t^{\prime}=t$, together with $h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+\textbf{c}\cdot\textbf{x}$, which ensures that the combination $\lambda+2\lambda_{1}$ does not renormalize; see also Ref. Jana et al. (2024). This further means $\chi+z=2$ exactly so long as nonlinearities remain relevant (in the scaling sense).
@@
In this Letter, we focus on the strong coupling phases of (2). We show that these strong coupling phases can be of two distinct kinds: (i) rough phase with positional short range order (SRO), but orientational long range order (LRO), or (ii) crumpled with both positional and orientational SRO. We extract the scaling exponents of the rough phase in $d\geq 2$, which show their nonuniversal nature: The variance $\Delta\equiv\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle\sim L^{2\chi(\gamma)}$, $10<\chi(\gamma)<$ 1 where $L$ is the linear size of the surface and $\overline{h}(t)$ is the average height at time $t$. Furthermore, the time-scale of relaxation $\tau(L)\sim L^{z(\gamma)}$. Both $\chi$ and $z$ are parametrized continuously by $\gamma\equiv\lambda_{1}/\lambda$, reflecting model parameter-dependent, nonuniversal scaling. In the crumpled phase, $\chi>1$.
@@
$g\equiv\frac{\lambda^{2}D}{\nu^{3}}k_{d},\,g_{1}\equiv\frac{\lambda_{1}^{2}D}{\nu^{3}}k_{d},\,g_{2}\equiv\text{sign}(\gamma)\sqrt{g_{1}}.$ (3)
Here $k_{d}=\frac{S_{d}}{(2\pi)^{d}}$, and $S_{d}$ is the surface area of a $d$-dimensional sphere. Considering $\lambda>0$ without any loss of generality, we notice that from (3) above that $g_{2}$ can be both positive and negative, depending upon the sign of $\gamma$. There are no one-loop corrections to $\lambda,\lambda_{1}$. The one-loop corrections to $\nu$ and $D$ can be obtained following Refs. Barabási and Stanley (1995); Forster et al. (1977); Jana et al. (2024), which allow us to calculate the one-loop RG flow equations for $g$, $g_{1}$ and hence $g_{2}$,
@@
where $\tilde{A}(\gamma)=\frac{9}{8}\gamma^{2}+\frac{11}{8}\gamma-\frac{1}{4}$. Considering 2D, the physically relevant dimension, we focus on the RG flow lines in the $g$-$g_{2}$ plane, with (0,0) as the only RG fixed point. Its stability properties are intriguing. It is stable (i.e., attractive) along the $g_{2}$-axis, but unstable (i.e., repulsive) along the $g$-axis. This indicates the existence of a separatrix, an invariant manifold under RG in the $g$-$g_{2}$ plane, that separates the stable phase from instability. While (0,0) is the only fixed point of both (22) and (23), it may be stable or unstable. Stability requires $dg/dl<0,dg_{2}/dl<0$, so that both $g(l),\,g_{2}(l)$ flow to zero in the long wavelength limit. Else when $dg/dl>0,dg_{2}/dl>0$, both $g(l),\,g_{2}(l)$ flow away from (0,0) giving instability. These two behaviors are demarcated by the separatrices $\tilde{A}(\gamma)=0$, giving
@@
for $\lambda_{1}>0$ and $\lambda_{1}<0$, respectively, as the separatrices, in $g$-$g_{2}$ plane passing through (0,0). The RG flow diagram in 2D is shown in Fig. 1(a). The red lines in Fig. 1(a) are the separatrices given in (6): the RG flow lines in the green regions bordered by the separatrices flow to the origin, giving the “weak coupling phase”. Although (0,0) is the stable fixed point in this region, the flow to (0,0) is so slow that $\nu$ and $D$ are infinitely renormalized, giving being super- or sub-logarithmically rough phases, as explored in Ref. Jana et al. (2024). In the remaining regions, both $g$ and $g_{2}$ grow monotonically with the “RG time” $l$. As soon as $g(l),\,g_{2}(l)\sim{\cal O}(1)$, which happens for a surface of finite size, the RG flow equations are outside the validity of the perturbative expansions. This is the perturbatively inaccessible “strong coupling phase”.
We first argue that the region confined between the two separatrices [red lines in Fig. 1(a)] actually consists of two subregions with distinct scaling properties (instead of just one region with a rough phase, as originally speculated in Ref. Jana et al. (2024)). To investigate that let us first consider the bare perturbation theory for $\nu$ in 2D, which is same as the RG calculations for $\nu$ except that now we extend the integrals over wave vectors down to an infrared cutoff $q_{\text{min}}=2\pi/L$. Setting $d=2$, we get for the effective $\nu_{e}$
@@
We thus find that for $\Bigl{(}\frac{1}{2}\gamma^{2}+\frac{5}{8}\gamma\Bigl{)}<0$ or $\Bigl{(}\frac{(g_{2})^{2}}{2g}+\frac{5g_{2}}{8\sqrt{g}}\Bigl{)}<0$ i.e. $-1.25<\gamma<0$, $\nu_{e}<\nu_{0}$. In fact, for sufficiently large $L/a$, $\nu_{e}<0$, giving divergence of $\langle({\bm{\nabla}}h)^{2}\rangle$ implying surface crumpling with positional and orientational SRO. In higher dimensions $d>2$, the one loop integrals are finite and hence the lower limit can be extended to 0, corresponding to the thermodynamic limit. Then for a range of $\gamma$, $-1.25<\gamma<0$ and low enough bare $\nu_{0}$, $\nu_{e}$ can turn negative, once again implying crumpling. We will revisit the region with crumpling below, shown in purple in Fig. 1(a), by using mode coupling methods, and come to a similar conclusion, giving us confidence about the existence of a crumpled phase. See Fig. 1(b) for a schematic phase diagram in the $\lambda-\lambda_{1}$ plane.
Figure 1: (a) RG flow diagram in the $g$-$g_{2}$ plane in 2D. Stable and unstable regions are divided by the separatrices (red lines). Stable (Unstable) regions are indicated by flow lines moving towards (away from) the origin. The arrows indicate the flow directions, delineating the stable and unstable regions. In this flow diagram (0,0) is the only fixed point that is unstable along the $g$-direction, but stable along the $g_{2}$-direction. (b) Phases in the $\lambda-\lambda_{1}$ plane. The green regions are the weak coupling logarithmically rough phase, the white regions are the algebraically rough phase, and the purple regions are the crumpled phase. In (a) and (b) the solid red line represents the locus of unstable fixed points where the roughening transition occurs. (c) Variation of $\chi$ in 2D in the phase space region where it is positive, with $\gamma$ as obtained from MCT. Central region where $\chi>1$, corresponds to a crumpled surface.
To study the unstable regions in the RG flow diagrams, we now set up an MCT calculation Bhattacharjee (1998); Basu and Frey (2009b) to extract the scaling behavior; see also Basu (2004); Basu and Bhattacharjee (2005) for applications of MCT in a slightly different context. The response $G({\bf k},\omega)$ and correlation $C({\bf k},\omega)$ functions have the following scaling form
$G({\bf k},\omega)=k^{-z}g\biggl{(}\frac{\omega}{k^{z}}\biggr{)}=-i\omega+\Sigma(k,\omega),$ (8)
$C({\bf k},\omega)=k^{-d-2\chi-z}f\biggl{(}\frac{\omega}{k^{z}}\biggr{)}.$ (9)
@@
Our argument is that the universal amplitude ratio $\Gamma^{2}/(D\lambda^{2})$ can be written down both from the one-loop diagrammatics of $G^{-1}({\bf k},\omega)$ and $C({\bf k},\omega)$. We use $\chi+z=2$, which is an exact result in the strong coupling phases (also consistent with the absence of vertex corrections at the one-loop order). Further assuming that $G^{-1}$ and $C$ are dominated by their respective one-loop contributions,which would hold if $z<2$ and $\chi>0$. Combining contributions from all the one-loop diagrams for $G({\bf k},\omega=0)$, we obtain sup
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\gamma^{2}\biggl{(}\frac{6\chi+4d-4}{d+2}\biggl{)}\biggl{]}.$ (11)
Next, considering the one-loop contributions to $C({\bf k},\omega=0)$, we find
@@
Now comparing RHS of Eqs. (58) and (71), we obtain a quadratic equation of $\chi$
@@
Here, $A=36\gamma^{2}+(6d+30)\gamma+(3d+6),\,B=(36d-54)\gamma^{2}+(8d^{2}+16d-42)\gamma+\biggl{(}\frac{d^{2}}{2}-4-d\biggl{)},\,C=(d-2)\gamma\bigl{[}\gamma(8d-8)+(2d^{2}+4d-6)\bigl{]}.$ Solving (13)
@@
On the separatrices (6), $B=0$ and $A=22.77$ corresponding to $\gamma=\gamma_{c1}=0.161$ and $A=19.69$ corresponding to $\gamma=\gamma_{c2}=-1.383$; $A$ changes from these boundary values within the unstable region between the separatrices. In particular, as $A$, depending upon $\gamma$, decreases, $\chi$ grows, eventually exceeding unity. At this point, our theory (2) breaks down, as $\langle({\bm{\nabla}}h)^{2}\rangle$ diverges with $L$, giving a crumpled surface. (We make a technical point here that as $\chi$ exceeds unity, higher order nonlinear terms not included in (2) become relevant (in a RG sense), making our theory given by Eq. (2) to breakdown.) Setting $\chi=1$ in (15), we get $\gamma=-1.043,-0.142$. Thus in the region $(-1.043<\gamma<-0.142)$ MCT predicts existence of a crumpled phase, which is the purple region in Fig. 1(a) and is slightly different from the region defined by $\nu_{e}=0$ (see above). Due to the breakdown of our theory, we cannot make any definitive conclusion about any scaling properties in the crumpled region.
Figure 2: Conjectured “Occam’s razor” global RG flows in the $g$-$g_{2}$ plane. Different speculated fixed points characterizing the rough and crumpled phases are marked. The broken lines are speculated fixed lines in the rough phase; see text.
[…]

### v3

@@ Abstract
We study a generalized Kardar-Parisi-Zhang (KPZ) equation [Jana et al., Phys. Rev. E 109, L032104 (2024)] that sets the paradigm for universality in roughening of growing nonequilibrium surfaces without any conservation laws but with competing local and nonlocal nonlinear effects. This equation in two dimensions exhibits two distinct strong coupling regimes: a rough phase and a crumpled phase, in addition to a weak coupling phase. The conformation fluctuations of such a rough surface are given by nonuniversal scaling exponents, with orientational long-range order and positional short-range order, whereas the crumpled phase has positional and orientational short-range order.
The Kardar-Parisi-Zhang (KPZ) equation Kardar et al. (1986); Krug (1997); Barabási and Stanley (1995), originally proposed as a model for growing surfaces without overhangs, provides a paradigm for nonequilibrium phase transitions. It undergoes a roughening transition between a smooth phase, statistically identical to a surface described by the linear Edwards-Wilkinson (EW) equation Edwards and Wilkinson (1982), and a perturbatively inaccessible strong coupling phase Barabási and Stanley (1995), in dimension $d>2$. Perturbative dynamic renormalization group (RG) Barabási and Stanley (1995); Täuber (2014) analysis has been successful in describing this roughening transition in an $\epsilon$ expansion, where $d=2+\epsilon$, with two dimensions (2D) being the lower critical dimension of the model. In $d\leq 2$, there exists only the strong coupling phase. A KPZ surface described by a single-valued height field $h({\bf x},t)$ measured with respect to an arbitrary base plane in the Monge gauge Nelson et al. (1989) displays universal scaling in the long wavelength limit, characterized by $z$ and $\chi$, the dynamic and roughness exponents Barabási and Stanley (1995). These are formally defined through the correlation function
@@
in the scaling limit, where $r=|{\bf x}|$ and $\phi$ is a dimensionless scaling function of its argument Barabási and Stanley (1995). In the smooth phase, $z=2$ and $\chi=1-d/2$ Barabási and Stanley (1995), and at the roughening transition at $d=2+\epsilon,\,\epsilon>0$, $z=2$ and $\chi=0$ Barabási and Stanley (1995); Täuber (2014); Janssen et al. (1999), which are accessible within perturbative RG calculations. In contrast, the strong coupling rough phase cannot be accessed within standard RG calculations Täuber (2014). Failure of RG to capture the rough phase has led to the development of alternative techniques. Notable are the self-consistent mode-coupling theories (MCT) Bouchaud and Cates (1993); Bhattacharjee (1998), which have predicted a variety of results on scaling within specific different calculational approaches. For example Ref. Hwa and Frey (1991) explored scaling at $d=1$, yielding $z=3/2$, a value known exactly due to the Galilean invariance and a fluctuation-dissipation theorem Barabási and Stanley (1995). Furthermore, Ref. Doherty et al. (1994) showed that $z$ rises from $z=3/2$ at $d=1$ to 222 around $d_{U}=3.6$, where $d_{U}$ is the upper critical dimension of the KPZ equation defined by the dimension in which $z$ touches the value 2 from a smaller numerical value at lower $d<d_{U}$. In contrast, Ref. Tu (1994) suggests $z<2$ at any finite dimension indicating $d_{U}=\infty$. Furthermore, confirming earlier functional RG results of Ref. Halpin-Healy (1990), Refs. Bhattacharjee (1998); Bhattacharjee and Bhattacharyya (2007); Basu and Frey (2004, 2009) used MCT to arrive, quite separately, at the same dimension-dependent exponents $\chi=(4-d)/6,z=(8+d)/6$, reasonably close to the more recent nonperturbative RG predictions Canet et al. (2010). This common functional RG and MCT finding suggests explicitly that $d_{U}=4$, supported independently by Ref. Fogedby (2005), using a third distinct technique. See also Ref. Colaiori and Moore (2001), which gives $d_{U}=4$ and the dynamic exponent $z$ in different dimensions. Furthermore, Ref. Canet and Moore (2007) illustrated the MCT solution for the KPZ equation at $d\geq 2$ that continues to $d_{U}=4$. More recent extensive numerical studies, however, appear to differ from the MCT predictions on the scaling exponents Kelling et al. (2017); Halpin-Healy (2013); Pagnani and Parisi (2015) and suggest $d_{U}>4$ Schwartz and Perlsman (2012); Pagnani and Parisi (2013); Kim and Kim (2013); Alves et al. (2014). See also Ref. Halpin-Healy (2025) for a very recent study on the scaling in the KPZ equation at $d>2$ by combining a variety of techniques. Furthermore, the results in Ref. Halpin-Healy and Takeuchi (2015) tend to indicate the absence of a finite $d_{U}$.
@@
The dynamics of a KPZ surface depends locally on surface fluctuations and hence cannot model nonequilibrium surfaces with nonlocal dynamics. Nonlocal effects, however, can be important in wide-ranging systems, including biological growth processes Santalla and Ferreira (2018), fast nonlocal transport Aharonov and Rothman (1996), and nonlocal stabilization of surfaces Krug and Meakin (1991); Nicoli et al. (2009). Recently, a generalized KPZ equation (hereafter G-KPZ equation) with nonlocal effects as a conceptual model with competing local and nonlocal nonlinear effects has been proposed Jana et al. (2024):
$\frac{\partial h}{\partial t}$ $=\nu\nabla^{2}h+\frac{\lambda}{2}(\bm{\nabla}h)^{2}$
$+$ $\quad\lambda_{1}\int d^{d}r^{\prime}Q_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))+\eta.$ (2)
@@
Here, $Q_{ij}({\bf r})$ denotes the longitudinal projection operator, which in Fourier space is expressed as $Q_{ij}({\bf k})=k_{i}k_{j}/k^{2}$, with ${\bf k}$ being a Fourier wave vector. Thus $Q_{ij}({\bf r})$, which is the inverse Fourier transform of $Q_{ij}({\bf k})$, is nonlocal in space. Physically, the term $\lambda_{1}\int d^{d}r^{\prime}Q_{ij}({\bf r-r^{\prime}})(\nabla_{i}h({\bf r^{\prime}},t)\nabla_{j}h({\bf r^{\prime}},t))$ contributes to the surface velocity normal to the base plane, $v_{p}=\partial h/\partial t$, and is nonlocal in ${\bm{\nabla}}h$. Furthermore, $\eta({\bf x},t)$ is a zero-mean Gaussian-distributed white noise with a variance $\langle\eta({\bf x},t)\eta({\bf 0},0)\rangle=2D\delta^{d}({\bf x})\delta(t).$ If $\lambda_{1}=0$, Eq. (2) reduces to the usual KPZ equation. Equation (2) serves as a minimal nonequilibrium model to study competition between local and nonlocal nonlinearities. For instance, as a consequence of these competitions, the G-KPZ equation (2) shows macroscopic properties dramatically different from the KPZ equation. It can have a stable, weak coupling sub- or superlogarithmically rough surface in 2D, characterized by nonuniversal scaling exponents. More strikingly and unlike the 2D KPZ equation, it also admits a roughening transition even in 2D, again characterized by nonuniversal exponents. Furthermore, at $d>2$ it shows a roughening transition different from that in the KPZ equation; see Ref. Jana et al. (2024), which principally focused on the weak coupling phase behaviors. Equation (2) admits a pseudo-Galilean invariance under the transformation $x_{i}^{\prime}=x_{i}-(\lambda+2\lambda_{1})c_{i}t$ and $t^{\prime}=t$, together with $h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+\textbf{c}\cdot\textbf{x}$, which ensures that the combination $\lambda+2\lambda_{1}$ does not renormalize; see Supplemental Material sup , see also Ref. Jana et al. (2024). This further means $\chi+z=2$ exactly, as long as nonlinearities remain relevant (in the scaling sense).
@@
In this Letter we focus on the strong coupling phases of (2). We show that these strong coupling phases can be of two distinct kinds: (i) a rough phase with positional short-range order (SRO) but orientational long-range order (LRO), or (ii) crumpled with both positional and orientational SRO. We extract the scaling exponents of the rough phase in $d\geq 2$. The variance $\Delta\equiv\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle\sim L^{2\chi(\gamma)}$, $0<\chi(\gamma)<1$, where $L$ is the linear size of the surface and $\overline{h}(t)$ is the average height at time $t$. Furthermore, the timescale of relaxation $\tau(L)\sim L^{z(\gamma)}$. Both $\chi$ and $z$ are parametrized continuously by $\gamma\equiv\lambda_{1}/\lambda$, reflecting model parameter-dependent, nonuniversal scaling. In the rough (crumpled) phase, $0<\chi<1\,(\chi>1)$.
Figure 1: (a) Phases and the RG flow lines in the $g$-$g_{2}$ plane at 2D. Stable and unstable regions are divided by the separatrices (red lines). Stable (unstable) regions are indicated by flow lines moving towards (away from) the origin. The arrows indicate the flow directions, delineating the stable and unstable regions. In this flow diagram (0,0) is the only fixed point, which is unstable along the $g$ direction but stable along the $g_{2}$ direction. (b) Phases in the $\lambda$-$\lambda_{1}$ plane. The green regions are the weak coupling logarithmically rough phase, the white regions are the algebraically rough phase, and the purple regions are the crumpled phase. In (a) and (b) the solid red line represents the locus of unstable fixed points where the roughening transition occurs. (c) Variation of $\chi$ in 2D in the phase space region where it is positive, with $\gamma$ as obtained from MCT. The central region where $\chi>1$ corresponds to a crumpled surface. See text.
@@
$g\equiv\frac{\lambda^{2}D}{\nu^{3}}k_{d},\,g_{1}\equiv\frac{\lambda_{1}^{2}D}{\nu^{3}}k_{d},\,g_{2}\equiv\text{sgn}(\gamma)\sqrt{g_{1}}.$ (3)
Here $k_{d}=\frac{S_{d}}{(2\pi)^{d}}$, and $S_{d}$ is the surface area of a $d$-dimensional unit sphere. Without any loss of generality, we are considering $\lambda>0$. We notice that from (3) above that $g_{2}$ can be both positive and negative, depending upon the sign of $\gamma$. There are no one-loop corrections to $\lambda,\lambda_{1}$. The one-loop corrections to $\nu$ and $D$ can be obtained following Refs. Barabási and Stanley (1995); Forster et al. (1977); Jana et al. (2024), which give the one-loop RG flow equations for $g$ and $g_{2}$:
@@
where $\tilde{A}(\gamma)=\frac{9}{8}\gamma^{2}+\frac{11}{8}\gamma-\frac{1}{4}$. Considering 2D, the physically relevant dimension, we focus on the RG flow lines in the $g$-$g_{2}$ plane, with (0,0) as the only RG fixed point. Its stability properties are intriguing. It is stable (i.e., attractive) along the $g_{2}$ axis but unstable (i.e., repulsive) along the $g$ axis. This indicates the existence of a separatrix, an invariant manifold under RG in the $g$-$g_{2}$ plane, that separates the stable phase from instability. While (0,0) is the only fixed point of both (4) and (5), it may be stable or unstable. Stability requires $dg/dl<0,dg_{2}/dl<0$, so that both $g(l),\,g_{2}(l)$ flow to zero in the long wavelength limit. Otherwise, when $dg/dl>0,dg_{2}/dl>0$, both $g(l),\,g_{2}(l)$ flow away from (0,0), giving instability. These two behaviors are demarcated by the separatrices $\tilde{A}(\gamma)=0$, giving
@@
for $\lambda_{1}>0$ and $\lambda_{1}<0$, respectively, as the separatrices, in a $g$-$g_{2}$ plane passing through (0,0). The RG flow diagram in 2D is shown in Fig. 1(a). The red lines in Fig. 1(a) are the separatrices given in (6). The RG flow lines in the green regions demarcated by the separatrices flow to the origin, giving the “weak coupling phase.” Although (0,0) is the stable fixed point in this region, the flow to (0,0) is so slow that $\nu$ and $D$ are infinitely renormalized, giving super- or sublogarithmically rough phases, as explored in Ref. Jana et al. (2024). In the remaining regions, both $g$ and $g_{2}$ grow monotonically with the “RG time” $l$. As soon as $g(l),\,g_{2}(l)\sim{\cal O}(1)$, which happens for a surface of finite size, the RG flow equations are outside the validity of the perturbative expansions. This is the perturbatively inaccessible “strong coupling phase.”
We first argue that the region confined between the two separatrices [red lines in Fig. 1(a)] actually consists of two subregions with distinct scaling properties (instead of just one strong coupling phase, as originally speculated in Ref. Jana et al. (2024)). To investigate that, first consider the bare perturbation theory for $\nu$ in 2D, which is same as the RG calculations for $\nu$ except that now we extend the integrals over wave vectors down to an infrared cutoff $q_{\text{min}}=2\pi/L$. Setting $d=2$, we get for the effective $\nu_{e}$,
@@
We thus find that for $\Bigl{(}\frac{1}{2}\gamma^{2}+\frac{5}{8}\gamma\Bigl{)}<0$ or $\Bigl{(}\frac{(g_{2})^{2}}{2g}+\frac{5g_{2}}{8\sqrt{g}}\Bigl{)}<0$, $-1.25<\gamma<0$, $\nu_{e}<\nu_{0}$. For a crumpled surface, $\langle\delta{\bf n}^{2}\rangle$ diverges for sufficiently large $L/a$, where $\delta{\bf n}$ is deviation of the local normal from a fully flat state. In the Monge gauge Nelson et al. (1989), to the lowest order in height fluctuations $\delta n\sim{\bm{\nabla}}h$. For sufficiently large $L/a$, $\nu_{e}<0$ gives divergence of $\langle({\bm{\nabla}}h)^{2}\rangle$, implying surface crumpling with positional and orientational SRO. In higher dimensions $d>2$, the one-loop integrals are finite and hence the lower limit can be extended to 0, corresponding to the thermodynamic limit. Then for a range of $\gamma$ given by $-1.25<\gamma<0$ and low enough bare $\nu_{0}$, $\nu_{e}$ can turn negative, once again indicating crumpling. We will revisit the region with crumpling below, shown in purple in Fig. 1(a), by using mode coupling methods and come to a similar conclusion, giving us confidence about the existence of a crumpled phase. See Fig. 1(b) for a schematic phase diagram in the $\lambda$-$\lambda_{1}$ plane.
To study the unstable regions in the RG flow diagrams, we now set up a one-loop MCT calculation Bhattacharjee (1998); Basu and Frey (2009) and extract the scaling behavior by using it; see also Chattopadhyay et al. (2000); Basu (2004); Basu and Bhattacharjee (2005) for applications of MCT in slightly different but related contexts. The response $G({\bf k},\omega)$ and correlation $C({\bf k},\omega)$ functions have the following scaling form:
$G^{-1}({\bf k},\omega)=k^{z}g\biggl{(}\frac{\omega}{k^{z}}\biggr{)}=-i\omega+\Sigma({\bf k},\omega),$ (8)
@@
Our argument is that the universal amplitude ratio $\Gamma^{2}/(D\lambda^{2})$ can be written down both from the one-loop diagrammatics of $G^{-1}({\bf k},\omega)$ and $C({\bf k},\omega)$. We use $\chi+z=2$, which is an exact result in the strong coupling phases (also consistent with the absence of vertex corrections at the one-loop order), further assuming that $G^{-1}$ and $C$ are dominated by their respective one-loop contributions,which would hold if $z<2$ and $\chi>0$. Combining contributions from all the one-loop diagrams for $G({\bf k},\omega=0)$, we obtain sup
@@
$+\gamma^{2}\biggl{(}\frac{6\chi+4d-4}{d+2}\biggl{)}\biggl{]}.$ (11)
From the one-loop contributions to $C({\bf k},\omega=0)$, we find
@@
Now comparing the right-hand side of Eqs. (11) and (12), we obtain a quadratic equation of $\chi$:
@@
Here, $A=36\gamma^{2}+(6d+30)\gamma+(3d+6),\,B=(36d-54)\gamma^{2}+(8d^{2}+16d-42)\gamma+\biggl{(}\frac{d^{2}}{2}-4-d\biggl{)},\,C=(d-2)\gamma\bigl{[}\gamma(8d-8)+(2d^{2}+4d-6)\bigl{]}.$ Solving (13),
@@
We take $\chi=\chi_{+}$, for with $\gamma=0$ only $\chi_{+}$ reduces to $\chi=1/2$ at $d=1$, the known exact value; incidentally, $\chi=\chi_{+}$ reduces to $\chi=1/3$ in 2D and $d_{U}=4$ obtained in an analogous MCT for the pure KPZ equation Bhattacharjee (1998) for $\gamma=0$. We now calculate $\chi$ in 2D. With $d=2$ in $A$, $B$, and $C$ defined above, we get $C=0$ and $\chi=\frac{-B+|B|}{2A}$, where $B=18\gamma^{2}+22\gamma-4$ and $A=36\gamma^{2}+42\gamma+12$. Validity of our MCT requires $\chi>0$. This means either $B<0$ and $A>0$, or $B>0$ and $A<0$. Notice that $B=0$ gives the separatrix (6) with $B<0$ corresponding to the unstable region in Fig. 1(a). We focus on the region $B<0$, which holds for $-1.383<\gamma<0.161$. Within this region,
On the separatrices (6), $B=0$ and $A=22.77$ corresponding to $\gamma=\gamma_{c1}=0.161$ and $A=19.69$, corresponding to $\gamma=\gamma_{c2}=-1.383$; $A$ changes from these boundary values within the unstable region between the separatrices. In particular, as $A$, depending upon $\gamma$, decreases, $\chi$ grows, eventually exceeding unity. This gives $\langle({\bm{\nabla}}h)^{2}\rangle$ diverging with $L$, giving a crumpled surface. [We make a technical point here that as $\chi$ exceeds unity, higher order nonlinear terms, e.g., $({\bm{\nabla}}h)^{2n},\,n>1$, not included in (2) should be relevant (in the scaling/RG sense). Thus Eq. (2), though it predicts the onset of the crumpled phase, cannot be used to study the crumpled phase.] Setting $\chi=1$ in (15), we get $\gamma=-1.043,-0.142$. Thus in the region $(-1.043<\gamma<-0.142)$ MCT predicts a crumpled phase, which is the purple region in Fig. 1(a) and is slightly different from the region defined by the condition $\nu_{e}=0$ [see Eq. (7) above]. Due to the breakdown of our theory, we cannot make any definitive conclusion about any scaling properties in the crumpled region.
Figure 2: Conjectured “Occam’s razor” global RG flows in the $g$-$g_{2}$ plane. Different speculated fixed points/lines characterizing the rough and crumpled phases are marked. The broken lines are speculated, fixed lines in the rough phase; see text.